### v1

# Don’t Break the Cache: An Evaluation of Prompt Caching for Long-Horizon Agentic Tasks

Abstract
Recent advancements in Large Language Model (LLM) agents have enabled complex multi-turn agentic tasks requiring extensive tool calling, where conversations can span dozens of API calls with increasingly large context windows. However, although major LLM providers offer prompt caching to reduce cost and latency, its benefits for agentic workloads remain underexplored in the research literature. To our knowledge, no prior work quantifies these cost savings or compares caching strategies for multi-turn agentic tasks. We present a comprehensive evaluation of prompt caching across three major LLM providers (OpenAI, Anthropic, and Google) and compare three caching strategies, including full context caching, system prompt only caching, and caching that excludes dynamic tool results. We evaluate on DeepResearchBench, a multi-turn agentic benchmark where agents autonomously execute real-world web search tool calls to answer complex research questions, measuring both API cost and time to first token (TTFT) across over 500 agent sessions with 10,000-token system prompts. Our results demonstrate that prompt caching reduces API costs by 45-80% and improves time to first token by 13-31% across providers. We find that strategic prompt cache block control, such as placing dynamic content at the end of the system prompt, avoiding dynamic traditional function calling, and excluding dynamic tool results, provides more consistent benefits than naive full-context caching, which can paradoxically increase latency. Our analysis reveals nuanced variations in caching behavior across providers, and we provide practical guidance for implementing prompt caching in production agentic systems.
Prompt Caching, Large Language Models, Agentic AI, Tool Calling, API Optimization
PricewaterhouseCoopers, U.S.
1 Introduction
Recent advancements in Large Language Model (LLM) agents have enabled complex, long-horizon agentic tasks that require extensive tool calling across multi-turn conversations (Ji, 2025). Through function calling, LLM agents can invoke external APIs, execute web searches, interact with databases, and perform domain-specific actions on behalf of users. As these agentic workloads grow in complexity, conversations can span dozens of API calls with context windows accumulating tens of thousands of tokens, leading to significant costs and latency overhead. To address this, major LLM providers including OpenAI, Anthropic, and Google now offer prompt caching, a feature that reuses previously computed key-value (KV) tensors from attention layers to avoid redundant computation on repeated prompt prefixes (OpenAI, 2026; Anthropic, 2026; Google Cloud, 2026a).
Figure 1: Prompt caching benefits (best cache mode per model). Percentage reduction in API cost and time to first token (TTFT) relative to a no-cache baseline. For each model, results correspond to the best-performing cache strategy. Asterisks denote statistically significant TTFT improvements ($p<0.05$).
While providers offer reduced pricing for cached input tokens, the benefits of prompt caching in real-world agentic workloads remain under-explored in the research literature. Existing work on KV cache optimization focuses primarily on inference-level memory management and compression (Kwon et al., 2023; Ge et al., 2023; Shi et al., 2024), rather than evaluating the enterprise-grade prompt caching features offered through provider APIs. Concurrent work has audited prompt caching across providers to detect timing side-channel vulnerabilities (Gu et al., 2025), but to our knowledge, no prior work has quantified the cost benefits of prompt caching or compared caching strategies for agentic workloads. This gap is particularly significant given the recent proliferation of long-running agents for deep research, coding assistance, and autonomous task completion, where prompt caching could substantially reduce operational costs and improve user experience through faster response times.
In this paper, we present the first comprehensive evaluation of prompt caching strategies for long-horizon agentic tasks across three major LLM providers (OpenAI, Anthropic, and Google) using four flagship models (Figure 1). We compare three caching strategies, including full context caching, system prompt only caching, and caching that excludes dynamic tool results. We evaluate on DeepResearchBench (Du et al., 2025), a multi-turn agentic benchmark where agents autonomously execute web search tool calls to answer complex research questions. Our evaluation spans over 500 agent sessions with 10,000-token system prompts, measuring both API cost and time to first token (TTFT) across all conditions.
Our evaluation reveals three key findings:
Prompt caching delivers substantial and consistent cost savings across all providers: All four models tested show statistically significant cost reductions when prompt caching is enabled. Cost savings range from 45% to 80% across providers. These savings are consistent across all three caching strategies, demonstrating that prompt caching provides reliable cost benefits regardless of the specific caching approach employed.
Latency improvements vary significantly across providers and require careful strategy selection: Time to first token improvements range from 13% to 31% across providers, though latency variance differs substantially between providers. Notably, the cache strategy that maximizes cost savings does not always maximize latency improvement, highlighting the importance of strategy selection based on optimization goals.
Strategic cache boundary control outperforms naive full-context caching: Providers abstract much of the caching mechanism, automatically triggering cache creation when token thresholds are exceeded. However, naively enabling full-context caching can paradoxically increase latency, as dynamic tool calls and results may trigger cache writes for content that will not be reused across sessions. By strategically controlling cache boundaries, such as caching only the system prompt or explicitly excluding tool results, practitioners can ensure that only stable, reusable content is cached. Our results show that system prompt only caching provides the most consistent benefits across both cost and latency dimensions.
2 Background
2.1 KV Cache and LLM Inference
Large Language Model inference consists of two distinct phases: the prefill phase, where the model processes the input prompt and generates attention key-value (KV) tensors, and the decode phase, where the model autoregressively generates output tokens (Pope et al., 2022). During prefill, the model computes attention over the entire input sequence, producing KV tensors that capture the contextual representations needed for subsequent generation. These KV tensors are stored in the KV cache and reused during decoding to avoid redundant computation, enabling efficient token-by-token generation (Not Lain, 2025).
As context windows have grown from thousands to millions of tokens, KV cache management has become a critical bottleneck in LLM serving (Shi et al., 2024). The memory footprint of KV caches scales linearly with sequence length and batch size, often consuming more GPU memory than the model weights themselves for long-context workloads. This has motivated extensive research on KV cache optimization, including memory management techniques such as PagedAttention (Kwon et al., 2023), which applies paging-style memory management to reduce fragmentation and waste, achieving 2-4x throughput improvements. Other approaches focus on KV cache compression through selective retention of important tokens (Ge et al., 2023), storage-compute tradeoffs that balance recomputation against cache loading (Jin et al., 2024), and shared prefix optimization for high-throughput inference (Juravsky et al., 2024; Wu and others, 2024; Sun et al., 2025; Zhou et al., 2024).
2.2 Prompt Caching in Provider APIs
While KV caching is a general inference optimization technique, prompt caching refers to the productized, provider-managed features that reuse KV tensors across API requests when prompts share common prefixes (OpenAI, 2026; Gim et al., 2024). By caching the KV tensors from the prefill phase, providers can skip redundant computation when subsequent requests begin with the same content, reducing both latency and cost for users.
Major LLM providers have implemented prompt caching with varying approaches. OpenAI offers automatic prompt caching on GPT-4o and newer models, where caching activates automatically for prompts exceeding a minimum token threshold, with cache hits occurring only for exact prefix matches (OpenAI, 2026, 2024a). Anthropic provides developer-controlled caching through explicit cache breakpoints, allowing users to specify which portions of their prompt should be cached, with configurable time-to-live (TTL) options (Anthropic, 2026, 2025c). Google offers both implicit caching, which activates automatically with no guaranteed cost savings, and explicit context caching, where developers create and reference caches with guaranteed discounts (Google Cloud, 2026a; Kilpatrick, 2025).
Implementation details such as minimum token thresholds (typically 1,024-4,096 tokens depending on model, see Table 4), TTL durations (ranging from 5 minutes to 24 hours), and pricing structures vary across providers and are subject to change (PromptHub, 2025; Microsoft Azure AI, 2025). These differences have practical implications for cache hit rates and cost optimization. Recent work has audited prompt caching across 17 providers, demonstrating that cache hits produce measurable TTFT reductions and identifying security vulnerabilities from timing side-channels (Gu et al., 2025). However, their focus on security auditing using synthetic prompts and smaller, older generation models does not compare caching strategies or evaluate cost and latency benefits for long-running agentic tasks on modern flagship models.
2.3 Agentic Workloads and Context Engineering
Recent advances in LLM agents have enabled complex, long-horizon tasks that extend far beyond single-turn question answering. Modern agentic applications including deep research assistants, coding agents such as Claude Code and Cursor, and autonomous task completion systems like Manus routinely execute 30-50 or more tool calls within a single session (Ji, 2025; Du et al., 2025; Mialon et al., 2023; Zhou et al., 2023; Drouin et al., 2024; Wei et al., 2025). Each tool call adds content to the conversation context, including the tool invocation, execution results, and the model’s subsequent reasoning, causing context windows to grow rapidly throughout the session.
This growth presents unique challenges for prompt caching (Guan et al., 2026; Laban et al., 2025). Recent work has proposed solutions for multi-turn caching scenarios (Jeong and Ahn, 2025; Yan et al., 2025). Unlike static question-answering scenarios where prompts are largely predetermined, agentic workloads feature dynamic, session-specific content that accumulates unpredictably. Tool results often contain user-specific data that will not benefit other sessions, and the interleaving of static system prompts with dynamic tool outputs complicates cache reuse. Context engineering strategies have emerged to manage these challenges, including treating external storage as extended memory and carefully structuring prompts to maximize cache efficiency (Ji, 2025; Lumer et al., 2025a). However, the effectiveness of prompt caching across different caching strategies in agentic workloads has not been comprehensively evaluated. Our work addresses this gap by measuring cost and latency benefits across controlled caching strategies on a multi-turn agentic benchmark.
3 Methodology
3.1 Experimental Setup
We evaluate prompt caching across three major LLM providers: OpenAI, Anthropic, and Google. For each provider, we select a flagship model: GPT-4o and GPT-5.2 from OpenAI, Claude Sonnet 4.5 from Anthropic, and Gemini 2.5 Pro from Google. These models represent the current state-of-the-art for agentic workloads and all support prompt caching through their respective APIs.
We use DeepResearchBench (Du et al., 2025) as our evaluation benchmark, a multi-turn agentic benchmark where agents autonomously execute web search tool calls to answer complex research questions. We selected this benchmark over alternatives such as other deep research benchmarks (FutureSearch: Bosse et al., 2025; Li et al., 2025) due to its focus on tool-intensive agentic workflows and real-world 100 PhD-level research tasks, each meticulously crafted by domain experts across 22 distinct fields. We implement our research agent using Deep Agents (LangChain, 2025), one of various open source libraries for creating long-running agents (Anthropic, 2025a; OpenAI, 2025; Google, 2025; Microsoft, 2023a, 2026, b; CrewAI, 2023; LlamaIndex, 2022; Hugging Face, 2024; OpenAI, 2024b; Agno, 2026). Each agent session begins with a research question and the agent iteratively calls a web search tool to gather information before synthesizing a comprehensive response. This benchmark reflects realistic agentic workloads where context windows grow dynamically through tool invocations and results.
For each model, we conduct 40 independent agent sessions per cache condition, with each session answering a unique research question from the benchmark. Sessions use a 10,000-token system prompt containing agent instructions for deep research, including guidance on tool usage, question decomposition, and report synthesis. Each session starts with a fresh context, ensuring that cache benefits are measured within individual multi-turn conversations rather than across sessions.
Figure 2: Prompt caching impact for normalized cost and time to first token (TTFT). Results use the system prompt only caching strategy. The no-cache baseline is normalized to 100% and lower values indicate better performance.
3.2 Latency Improvement
3.3 Cache Mode Implementation
We implement four cache conditions to systematically evaluate prompt caching strategies (see Appendix C, Figures 58 for visual illustrations). To control cache boundaries precisely, we use unique identifiers (UUIDs) to break the cache at specific points in the prompt, ensuring that content after the UUID is not cached from previous requests.
No Cache (Baseline): A UUID is prepended to the beginning of the system prompt, breaking the cache immediately and forcing the model to recompute all tokens. This serves as our baseline condition where no caching benefits are realized. In real-world agentic tasks, this symbolizes including dynamic content, such as timestamps and user information, to the system prompt on inference time (Ji, 2025).
Full Context Caching: No UUIDs are added, allowing the provider’s caching mechanism to operate automatically. OpenAI and Google enable prompt caching automatically for eligible requests, while Anthropic requires explicit cache breakpoints in the API request. This condition represents naive caching where practitioners enable the feature without additional optimization.
System Prompt Only Caching: A UUID is appended to the end of the system prompt, breaking the cache at this boundary. This ensures that only the static system prompt is cached, while the dynamic conversation history, tool calls, and tool results are recomputed on each request.
Exclude Tool Results Caching: UUIDs are appended both after the system prompt and after each tool result. This strategy ensures that tool results, which are dynamic and session-specific, do not contribute to the cache. We found this dual-UUID approach necessary because provider-level KV cache handling can vary, and explicit boundaries provide more predictable caching behavior.
3.4 Evaluation Protocol
We measure two primary metrics across all conditions: API cost and time to first token (TTFT).
Cost: We calculate cost using token counts reported in API responses, distinguishing between standard input tokens, cached input tokens (cache reads), and cache creation tokens (cache writes). Each token type is multiplied by the corresponding provider pricing (see Appendix A, Table 3) to compute total cost per session. Cost is aggregated across all API calls within a session.
Time to first token (TTFT): We measure TTFT using streaming responses, recording the time from request initiation to receipt of the first response chunk. TTFT captures the latency improvement from skipping prefill computation on cached tokens, making it the most relevant latency metric for prompt caching evaluation.
Prior to each experimental condition, we execute warmup calls to prime the cache and record cache creation tokens separately from evaluation runs. Between conditions for different cache modes, we wait sufficient time (exceeding 24 hours) to ensure cache entries expire based on provider TTL policies, preventing cross-condition cache contamination.
3.5 Statistical Analysis
We compare each cache condition against the no-cache baseline using independent samples t-tests. Statistical significance is determined at $\alpha=0.05$. For each model and cache mode, we report mean cost, mean TTFT, percentage improvement over baseline, and p-values. Sample sizes are $n=40$ per condition for all models.
4 Results
4.1 Overall Results
Table 1 summarizes the prompt caching benefits across all four models using the best-performing cache mode for each model. All experiments show statistically significant improvements ($p<0.05$). Cost savings range from 41% to 80% across models, while time to first token improvements range from 6% to 31%.
Table 1: Prompt caching benefits by model using the best-performing cache mode for each. Cost savings and TTFT improvement are relative to the no-cache baseline. Bold indicates highest value per metric.
Model Cache Mode Cost $\downarrow$ TTFT $\downarrow$
OpenAI GPT-5.2 Excl. Tool Results 79.6% 13.0%
Claude Sonnet 4.5 System Prompt 78.5% 22.9%
Gemini 2.5 Pro System Prompt 41.4% 6.1%
OpenAI GPT-4o System Prompt 45.9% 30.9%
4.2 Cost Reduction
Table 2: Full comparison of cache modes across all models. Cost savings and TTFT improvement are relative to the no-cache baseline. Negative TTFT values indicate regression.
Model Cache Mode Cost $\downarrow$ TTFT $\downarrow$
OpenAI GPT-5.2 No Cache (Baseline)
Full Context 79.3% 9.5%
System Prompt 81.4% 10.5%
Excl. Tool Results 79.6% 13.0%
Claude Sonnet 4.5 No Cache (Baseline)
Full Context 77.8% 21.8%
System Prompt 78.5% 22.9%
Excl. Tool Results 78.1% 20.9%
Gemini 2.5 Pro No Cache (Baseline)
Full Context 38.3% 6.0%
System Prompt 41.4% 6.1%
Excl. Tool Results 27.8% -2.9%
OpenAI GPT-4o No Cache (Baseline)
Full Context 47.8% -8.8%
System Prompt 45.9% 30.9%
Excl. Tool Results 46.8% 28.1%
(a) Normalized cost distribution by cache mode
(b) Normalized TTFT distribution by cache mode
Figure 3: Normalized cost and time to first token (TTFT) distributions by model and cache strategy. The no-cache baseline is normalized to 100% and lower values indicate better performance. Cost reductions are consistent across cache strategies, while TTFT improvements vary significantly, with full-context caching sometimes underperforming more selective strategies.
Prompt caching delivers substantial cost reductions across all providers and cache strategies. As shown in Table 2 and Figure 2, all cache modes achieve cost savings compared to the no-cache baseline for all four models. Cost reductions range from 79-81% for GPT-5.2, 78-79% for Claude Sonnet 4.5, 46-48% for GPT-4o, and 28-41% for Gemini 2.5 Pro depending on the cache mode selected. The consistency of cost savings across cache strategies suggests that the primary driver of cost reduction is caching the large system prompt, which remains stable across all requests within a session. Additional caching of conversation history and tool calls provides marginal incremental benefit for cost, as these components are smaller relative to the system prompt in our experimental setup.
Time to first token improvements show greater variation across providers compared to cost savings. GPT-4o shows 28-31% improvement with system prompt only and exclude tool results strategies, while full context caching exhibits a slight regression of 8.8%, suggesting that caching dynamic content can introduce overhead that negates latency benefits. Claude Sonnet 4.5 demonstrates consistent TTFT improvements across all cache strategies, ranging from 20.9% to 22.9%, including full context caching. This indicates that provider implementations differ in how they handle dynamic content caching. GPT-5.2 shows 13.0% improvement with the exclude tool results strategy, while Gemini 2.5 Pro shows 6.1% improvement with system prompt only caching. TTFT measurements exhibit natural variance due to factors including server load, network conditions, and provider infrastructure. This variance is reflected in the box plot distributions in Figure 3.
4.3 Cache Strategy Comparison
Figure 3 presents normalized cost and TTFT distributions across all four cache strategies for each model. The results reveal important differences in how cache strategies perform across providers.
For cost optimization, all three caching strategies (full context, system prompt only, and exclude tool results) provide similar benefits within each model. The differences between strategies are typically within 2-4 percentage points, indicating that the system prompt, which is cached in all strategies, drives the majority of cost savings.
For latency optimization, the choice of cache strategy has a more pronounced impact. System prompt only caching and exclude tool results caching consistently outperform full context caching for TTFT improvement. For some models, full context caching shows no improvement or slight regression, while other strategies achieve 28-31% improvement. The likely explanation is that full context caching triggers cache writes for dynamic tool calls and results, introducing overhead that offsets the benefits of cache reads.
5 Discussion
5.1 Strategic Cache Boundary Control
Our results demonstrate that strategic control over cache boundaries is essential for maximizing prompt caching benefits in agentic workloads. The key insight is that providers abstract much of the underlying caching mechanism, automatically triggering cache creation when token thresholds are exceeded. Without explicit boundary control, this automatic behavior can cache dynamic, session-specific content that will not be reused, leading to cache write overhead without corresponding read benefits.
The most effective strategy is to ensure that only stable, reusable content is cached. In agentic applications, the system prompt is the most stable component, containing agent instructions, tool definitions, and persona guidelines that remain constant across sessions. Conversation history, tool calls, and tool results are inherently dynamic and session-specific, making them poor candidates for cross-session caching. Practitioners should avoid including dynamic values in the system prompt itself. Common patterns that inadvertently break the cache include timestamps, datetime strings, session identifiers, or user-specific information embedded in the system prompt. If such dynamic information is necessary, it should be placed at the end of the system prompt to maximize the cacheable prefix. This ensures that the majority of the system prompt benefits from cache hits while only the dynamic suffix requires recomputation (Ji, 2025).
Similarly, dynamic function calling can break the cache when tool definitions change between requests. Modern agentic systems increasingly leverage dynamic tool discovery and registration through protocols such as the Model Context Protocol (MCP) (Model Context Protocol, 2026), where available tools may vary based on connected servers or runtime context (Lumer et al., 2025a). When tool definitions are included in the prompt, any change to the available tool set invalidates the cached prefix. A practical strategy is to maintain a fixed set of general-purpose, reusable functions (such as code execution, file operations, and bash/shell commands), while implementing dynamic capabilities through code generation rather than traditional function calling (Ji, 2025; Wang et al., 2024; Jones and Kelly, 2025; Anthropic, 2025b; Varda and Pai, 2025; Hacker News, 2025). Those prior methods of dynamic function calling, while achieving strong retrieval and execution accuracy, can prevent prompt caching usage (Lumer et al., 2024, 2025b; Chen et al., 2024; Zheng et al., 2024; Chen et al., 2024; Wu et al., 2024).
5.2 Tool Call Caching Considerations
For long-running agentic sessions with 30-50 or more tool calls, practitioners may consider caching tool calls and results to further reduce costs. However, this approach involves tradeoffs. Cache creation incurs a cost and latency overhead on the first request, which is only amortized if subsequent requests benefit from cache reads. For tool calls that produce highly variable results or that are unlikely to be repeated, caching provides no benefit and may introduce unnecessary overhead.
Furthermore, common context management strategies in agentic systems can interact poorly with tool call caching. Techniques such as summarizing or pruning old tool calls to manage context length (Ji, 2025) inherently modify the conversation history, breaking any cached representations of that content. If an application employs such strategies, caching tool calls becomes counterproductive. The emerging pattern for agentic applications is to maintain a large, stable system prompt that benefits from caching, while treating tool calls and results as dynamic content that may be summarized, pruned, or otherwise managed throughout the session.
5.3 Provider Implementation Variability
Provider implementations of prompt caching differ in important ways that affect practical deployment. Minimum token thresholds for cache eligibility range from 1,024 to 4,096 tokens depending on the provider and model (see Appendix A). Time-to-live (TTL) durations vary from 5 minutes to 24 hours, affecting whether cached content remains available across user sessions. Some providers offer automatic caching that activates without developer intervention, while others require explicit API parameters to enable caching. Enterprise deployments may also leverage dedicated caching infrastructure (Liu et al., 2024; Cheng et al., 2024; Yao et al., 2025; Cheng et al., 2025) to further optimize performance. These implementation details are subject to change and practitioners should consult current provider documentation when designing caching strategies.
Our results also reflect natural variance in API response times due to factors including server load, geographic distribution, and infrastructure differences across providers. When evaluating prompt caching benefits, practitioners should conduct experiments representative of their specific workloads and usage patterns rather than relying solely on published benchmarks. Practitioners should also be aware of security considerations, as recent work has demonstrated that prompt caching can introduce timing side-channels that may leak information about cached content (Wu et al., 2025; Gu et al., 2025).
6 Conclusion
Recent advancements in Large Language Model (LLM) agents have enabled complex multi-turn agentic tasks requiring extensive tool calling, where conversations can span dozens of API calls with increasingly large context windows. While major LLM providers offer prompt caching to reduce costs and latency, the benefits of these features remain under-explored in the research literature. In this work, we present a comprehensive evaluation of prompt caching across three major LLM providers (OpenAI, Anthropic, and Google) using four flagship models. We compare three caching strategies, including full context caching, system prompt only caching, and caching that excludes dynamic tool results. We evaluate on DeepResearchBench, a multi-turn agentic benchmark where agents autonomously execute web search tool calls to answer complex research questions, measuring both API cost and time to first token (TTFT) across over 500 agent sessions with 10,000-token system prompts. Our results demonstrate that prompt caching reduces API costs by 45-80% and improves time to first token by 13-31% across providers. We find that strategic cache boundary control, such as excluding dynamic tool results, provides more consistent benefits than naive full-context caching, which can paradoxically increase latency. As agentic systems continue to trend toward longer-running sessions with dozens of tool calls, strategic prompt caching becomes critical for reducing operational costs and improving user experience, and our findings provide practitioners with guidance for implementing prompt caching in production agentic systems.
Appendix A Prompt caching pricing at time of evaluation
Table 3: Token pricing as of early January 2026, used in our cost analysis (USD per 1M tokens unless otherwise noted). “Cached input” refers to cache-hit tokens. “Cache write” denotes cache creation costs when explicitly priced. Google additionally charges for context cache storage (\$4.50 per million tokens per hour), which is accounted for separately in our analysis. Pricing reflects public provider documentation at the time of evaluation. (OpenAI, 2026; Anthropic, 2026; Google Cloud, 2026b)
Provider Model Input Output Cached Write
OpenAI GPT-4o 2.50 10.00 1.25
OpenAI GPT-5.2 1.75 14.00 0.175
Anthropic Claude Sonnet 4.5 3.00 15.00 0.30 3.75
Google Gemini 2.5 Pro ($\leq$200K) 1.25 10.00 0.125
Google Gemini 2.5 Pro (>>200K) 2.50 15.00 0.250
Table 4: Minimum prompt length (in tokens) required for prompt caching to apply, as of early January 2026. Prompts shorter than these thresholds cannot benefit from caching, even when caching features are enabled. Thresholds reflect public provider documentation at the time of evaluation. (OpenAI, 2026; Anthropic, 2026; Google Cloud, 2026a)
Provider Model Min. Tokens
OpenAI GPT-4o 1,024
OpenAI GPT-5.2 1,024
Anthropic Claude Sonnet 4.5 1,024
Google Gemini 2.5 Pro 4,096
Appendix B Prompt Caching Mechanism
Figure 4 illustrates the fundamental mechanism underlying prompt caching. When a request is processed, the system checks whether the prompt prefix matches previously cached content. A cache hit occurs when the entire prefix matches exactly, allowing the system to reuse previously computed KV tensors (shown in green). A cache miss occurs when any token differs from the cached content, even at the very beginning (shown with an orange indicator), forcing complete recomputation of all tokens (shown in gray).
[…]
Figure 4: Prompt caching requires exact prefix matches. Different shades represent message types in agentic conversations: brightest (system prompt), light gray (human messages), medium gray (AI messages), darker gray (tool calls), and darkest (tool results). Cache hit: The prompt prefix matches a previously seen request exactly, so cached KV tensors are reused (green) and only new tokens appended at the end require computation (gray). Cache miss: Any difference in the prefix—even a single token at the beginning (orange)—prevents cache reuse, forcing full recomputation of all tokens (gray).
Appendix C Cache Strategy Implementations
The following figures illustrate the four cache strategies evaluated in this work. Since prompt caching operates on exact prefix matches, we use UUIDs (indicated by red bars) to control cache boundaries. Static content placed before the UUID forms the cacheable prefix; content after the UUID varies between requests and prevents prefix matches beyond that point.
C.1 No Cache (Baseline)
UUIDSystem PromptHuman + AITool Call + ResultContinued…UUID (cache breaker)Not cached (recomputed)
Figure 5: No Cache (Baseline): A unique UUID prepended to the start of the system prompt ensures no prefix match is possible with any prior request, forcing full recomputation of all tokens every time.
C.2 Full Context Caching
System PromptHuman + AITool Call + ResultContinued…No UUIDAutomatically cached by provider
Figure 6: Full Context Caching: No UUIDs are added, allowing the provider to automatically cache the entire prompt prefix. However, this may cache dynamic content (e.g., tool results) that varies between sessions, potentially triggering cache writes without corresponding cache hits.
C.3 System Prompt Only Caching
System PromptUUIDHuman + AITool Call + ResultContinued…UUIDCachedNot cached
Figure 7: System Prompt Only Caching: A UUID appended after the system prompt breaks the cacheable prefix at this boundary. The static system prompt (placed at the beginning) benefits from prefix caching, while dynamic conversation content (placed after) is recomputed each request.
C.4 Exclude Tool Results Caching
System PromptUUIDHuman + AITool CallTool ResultUUIDNext Human + AITool CallTool ResultUUIDUUID (after system + each tool result)CachedNot cached
Figure 8: Exclude Tool Results Caching: UUIDs are appended after the system prompt and after each tool result to break the cacheable prefix at these boundaries. This prevents session-specific tool results from being cached, avoiding cache writes for content unlikely to produce future cache hits. Furthermore, this mirrors cache-breaking context engineering strategies that prune or summarize past tool calls.